# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v2

@@ Appendix H Inference Details
Appendix I Pruning Details
I.1 Offline One-shot Pruning
@@
Figure 8: Pruning Dependency in the Qwen2.5
@@
To ensure hardware-efficient matrix multiplication during inference, we set the pruning granularity to 64 for embedding dimension and 256 for the intermediate dimension. This approach results in multiple width-pruned model configurations, as show in Table 8. We calculate channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
Table 8: Perplexity (PPL) on WikiText-2 for different model architectures with 20% and 30% width pruning.
@@
I.2 Recovery Training in Pruning
@@
Table 9: Performance comparison of Qwen2.5 models after pruning. We report 5-shot performance on mmlu and cmmlu, 4-shot for gsm8k, zero-shot for ARC-Challenge, HellaSwag. * denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
@@
Performance across Knowledge Assessment & Reasoning Tasks
Table 9 summarizes the performance of pruned models across various downstream tasks, including MMLU and CMMLU for knowledge assessment, alongside ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation. Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks. At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%. When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance. In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
@@
Table 10: Results with different pruning strategies on domain specific tasks.
@@
Appendix J FLOPs Calculation
@@
Appendix K Extensive Attempts on Generalization Enhancement via RL
Table 11: Full results during the RL exploration
@@
While attempting to apply reinforcement learning (RL), we conduct extensive experiments, with results shown in Appendix K. Initially, we find that directly training a 7B model with RL yields excellent results, surpassing the LLM Prototype and RFT-trained models. Subsequently, we attempt to train smaller models (3B and 0.5B) and discover that RL is ineffective on the 0.5B model, because the model tends to generate shorter or even meaningless Chain-of-Thought (CoT), leading to degraded output diversity as well as that without CoT. To address the issue above, we explore another approach, transferring out-of-domain capabilities into the 0.5B model via Knowledge Distillation (KD). However, we observe that the model struggles with format errors due to underfitting. Ultimately, after applying RL training on this model, we notice an impressive improvement on model performance, with both an absolute 1% AR increase on in-domain/out-of-domain datasets compared to the 0.5B RFT-trained model.

### v3

@@ Appendix G Inference Details
Appendix H Pruning Details
H.1 Offline One-shot Pruning
@@
Figure 8: Width pruning dependency in the Qwen2.5 model.
@@
To ensure hardware-efficient matrix multiplication during inference, we set the pruning granularity to 64 for embedding dimension and 256 for the intermediate dimension. This approach results in multiple width-pruned model configurations, as show in Table 7. We calculate channel importance scores $CI^{i}_{\text{embedding}}$ and $CI^{j}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 7 presents the optimal configurations for width pruning at ratio of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
Table 7: Perplexity(PPL) evaluation on WikiText-2 comparing model architectures with one-shot width pruning at 20% and 30% pruning ratios.
@@
H.2 Recovery Training in Pruning
@@
Performance across Knowledge Assessment & Reasoning Tasks
Table 8 summarizes the performance of pruned models across various downstream tasks, including MMLU[79] and CMMLU[80] for knowledge assessment, alongside ARC-Challenge[81] and HellaSwag[82] for commonsense reasoning evaluation, and GSM8k[83] for logical reasoning. Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks. At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%. When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance. In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
Table 8: Performance comparison of Qwen2.5 models after pruning. We report 5-shot performance on MMLU and CMMLU, 4-shot for GSM8k, zero-shot for ARC-Challenge, HellaSwag. * denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction fine-tuning.
@@
Table 9: Results in different pruning strategies on domain-specific tasks.
@@
Appendix I FLOPs Calculation
@@
Appendix J Extensive Attempts on Generalization Enhancement
Table 10: In-domain and out-of-domain performance of Qwen2.5 models using different knowledge transfer strategies.
@@
Knowledge Transfer using Qwen2.5-3B (RL)
Qwen2.5-0.5B (RFT*) 93.82% 89.13%
Qwen2.5-0.5B (RFT*+RL) 95.72% 90.09%
@@
Knowledge Transfer using Qwen2.5-7B (RL)
Qwen2.5-0.5B (RFT*) 95.29% 90.28%
Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
Qwen2.5-0.5B (KD*) 95.32% 90.59%
Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
Figure 9: Response lengths of different models during RL training
Table 11: Response lengths of different models, represented as mean $\pm$ standard deviation.
Model Response Length
QwQ-32B (ICL) 316.8 $\pm$ 111.4
Qwen2.5-7B (RL) 86.4 $\pm$ 12.9
Qwen2.5-3B (RL) 61.7 $\pm$ 9.4
Qwen2.5-0.5B (KD*) 86.3 $\pm$ 12.6
Qwen2.5-0.5B (KD*+RL) 85.3 $\pm$ 10.8
Upon discovering that using RFT or KD alone leads to significant degradation in out-of-domain performance for super-tiny models, we conduct extensive experiments to enhance generalization capabilities, as illustrated in Appendix J. Initially, we prioritize Reinforcement Learning (RL) for this task. As expected, 7B RL-trained models achieve excellent results, surpassing the LLM Prototype and RFT-trained models. Subsequently, we attempt direct RL training on smaller models, such as 3B and 0.5B. The results are less favorable this time. While 3B RL-trained models exhibit considerable generalization capabilities, 0.5B RL-trained models show no improvement in generalization, underperforming even compared to 0.5B RFT-trained models. Through observations and analyses, we conclude that this is because models of such small size tend to generate low-quality Chain-of-Thought (CoT), as shown in Figure 9, leading to degraded output diversity, akin to scenarios where CoT is not used.
To address the challenge of 0.5B models’ inabilities to generate high-quality CoT, we propose a hybrid strategy. Given the favorable results from 7B RL-trained models, we decide to use 7B RL-trained models as our starting point. Then, we transfer out-of-domain capabilities and the abilities of generating high-quality CoT from 7B RL-trained models to 0.5B models via Rejection sampling Fine-Tuning (RFT) or knowledge distillation (KD), which are denoted as RFT* and KD*, respectively. Experimental results indicate that this knowledge transfer indeed outperforms direct RL training on 0.5B models in enhancing generalization and preserving the abilities of generating high-quality CoT.
So far, 0.5B KD* models posses the abilities to generate high-quality CoT, but in terms of out-of-domain Achievable Rate (AR), there is no considerable improvement compared to 0.5B RFT-trained models. Consequently, we conduct another round of RL training on 0.5B KD* models with data generated by the LLM prototype for enhancing generalization, similar to how we initially improve 7B models’ generalization with RL. Ultimately, experimental results present that after the second RL training, the 0.5B super-tiny models exhibit impressive improvement on model performance, with both an absolute increase of over 1% AR increase on in-domain and out-of-domain datasets compared to 0.5B RFT-trained models.
Table 12: Ablation of knowledge transfer strategies on Qwen2.5-0.5B models.
Model AR
AR
(Out-of-Domain)
Qwen2.5-0.5B (RFT*) 95.29% 90.28%
Qwen2.5-0.5B (KD*[FKL]) 95.08% 90.19%
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 95.32% 90.59%
Qwen2.5-0.5B (KD*[AKL]) 93.93% 88.37%
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 95.82% 90.85%
Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
Qwen2.5-0.5B (KD*[FKL]+RL) 95.66% 92.21%
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 95.99% 92.34%
Qwen2.5-0.5B (KD*[AKL]+RL) 95.85% 92.10%
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 96.10% 92.03%
Ablation of Different Teacher Models
As illustrated in Appendix J, a teacher model with inherently superior capabilities is crucial for enhancing student model performance, particularly regarding out-of-domain capabilities. A better teacher model not only leads to better performance but also ensures greater stability in the student models’ results, driven by the generation of higher-quality CoT.
Response Lengths of Different Reasoning Models
As can be seen in Table 11, the average response lengths of our models are more than three times shorter than that of QwQ-32B. For direct RL training, longer response lengths appear to yield better results. However, this conclusion is not entirely valid. For instance, in the case of KD* models, there is almost no change in response lengths before and after RL, yet there is a significant difference in out-of-domain performance.
Ablation of Knowledge Transfer Strategies
During the knowledge transfer stage from 7B RL-trained Qwen2.5 models, we investigate the effectiveness of different knowledge transfer strategies, as illustrated in Appendix J. Apart from the KD*[AKL] strategy, which encounters a certain degree of underfitting, other experimental results demonstrate the robustness of different knowledge transfer strategies. Specifically, utilizing a loss with Cross Entropy (CE) is preferable to not incorporating CE during KD, as the high-quality CoT generated by 7B RL-trained models facilitates faster convergence with CE loss. Conversely, when the teacher model generates low-quality CoT, using CE alone, equivalent to RFT, can lead to overfitting, as evidenced in Appendix J, where the 3B RL-trained model as the teacher model exhibits poorer performance.
Appendix K Experiments on Public Datasets
Table 13: Performance of different models on BFCL and xlam-function-calling-60k datasets.
BFCL In-domain
Model Overall Non-live Live Acc
Qwen2.5-0.5B 55.14 62.29 47.98 55.88
Hammer2.0-0.5B [85] 59.31 67.00 51.62 /
Qwen2.5-72B 83.55 88.13 78.98 79.47
Qwen2.5-7B (RL) 81.85 86.27 77.42 78.55
Qwen2.5-0.5B (RL) 55.15 64.48 45.82 65.68
Qwen2.5-0.5B (RFT) 56.64 64.50 48.78 74.08
Qwen2.5-0.4B-Depth (RFT) 42.27 54.56 29.98 71.55
Qwen2.5-0.4B-Width (RFT) 42.60 56.17 29.02 72.98
Qwen2.5-0.5B-W8A16 (RFT) 56.42 64.42 48.41 74.38
Qwen2.5-0.5B-W4A16 (RFT) 49.54 58.52 40.56 72.23
Qwen2.5-0.5B-W8A8 (RFT) 55.68 62.88 48.48 74.06
Knowledge Transfer using Qwen2.5-7B (RL)
Qwen2.5-0.5B (KD*[FKL]) 59.44 69.06 49.81 72.97
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 54.28 62.00 46.56 66.55
Qwen2.5-0.5B (KD*[AKL]) 58.45 67.15 49.74 71.90
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 58.97 68.12 49.81 71.70
Qwen2.5-0.5B (KD*[FKL]+RL) 60.42 71.02 49.81 75.87
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 59.84 70.00 49.67 77.02
Qwen2.5-0.5B (KD*[AKL]+RL) 60.50 69.56 51.44 76.38
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 60.79 69.46 52.11 76.68
To validate the effectiveness of our proposed method, we conduct experiments on publicly available datasets. We randomly partition 90% of the xlam-function-calling-60k dataset [69] for training, with the remaining 10% serving as an in-domain test set. For the training set, instead of utilizing the provided labels, we employ Qwen2.5-72B-Instruct [11] to generate reference labels. Furthermore, we design a judge based on QwQ-32B [44] to assess the accuracy of the labels generated by the LLM. This process can be viewed as prototyping. Given that function calling is a standardized task, we use the official function call prompt template instead of developing a specialized agent for answer generation. The design of our judge is inspired by the Format Checker and Semantic Checker proposed in [69]. Through this procedure, we ultimately curate 50,201 valid query-tools-answers pairs from the initial 54k training samples for subsequent training.
Due to the complex function call formats within this dataset, we redesign the reward mechanism for Reinforcement Learning (RL). The reward is defined as:
$Reward=\alpha\cdot Reward_{format}+\beta\cdot Reward_{answer}$ (11)
where $Reward_{format}$ is 1 if and only if all the following conditions are met, and 0 otherwise: (1) The model’s output contains exactly one "<think>…</think>" block; (2) The function call within the answer can be parsed by the official Qwen prompt template and can be loaded as a JSON object; (3) The function name in the answer exists within the tools provided in the prompt, and the parameter types conform to their definitions, which is implemented using Pydantic[86].
For $Reward_{answer}$, we first check whether the number of functions in the model’s output is equivalent to that of the ground truth: If not, $Reward_{answer}$ is 0; Otherwise, $Reward_{answer}$ is defined as follows:
$Reward_{i}=\begin{cases}\frac{\lvert P_{i,pred}\cap P_{i,label}\rvert}{\lvert P_{i,pred}\cup P_{i,label}\rvert},&\text{if }\lvert P_{i,pred}\cup P_{i,label}\rvert>0\\ 1,&\text{otherwise}\end{cases}$ (12)
$Reward_{answer}=\frac{1}{N}\sum Reward_{i}$ (13)
where $N$ is the number of function calls, and $P_{i}$ denotes the set of parameters for the $i$-th function call. During training, we set $\alpha=1$ and $\beta=2$. The baseline variant of Reinforce++ [50] is selected as our reinforcement learning method.
We evaluate the trained model on both in-domain and out-of-domain dataset. For in-domain evaluation, we utilize the aforementioned test set with official answers as the ground truth. A test instance is considered correct only if the model’s output exactly matches the answer (i.e., identical function names and key-value pairs for arguments); all other cases are deemed incorrect. Accuracy is calculated based on this criterion. Additionally, we employ the Berkeley Function Call Leaderboard (BFCL)[70] for out-of-domain evaluation. The final experimental results are presented in Table 13.
The experimental results indicate that the 0.5B trained model achieves in-domain accuracy comparable to the Qwen2.5-72B model. Furthermore, it demonstrates a substantial improvement over Qwen2.5-0.5B-Instruct in out-of-domain scenarios. We also compare our method with Hammer2.0-0.5B[85], the previous SOTA 0.5B model, which is also fine-tuned from Qwen2.5-0.5B-Instruct8. Our approach achieves superior performance, demonstrating the effectiveness of our proposed method. In terms of quantization, experimental results demonstrate that applying W8A8 and W8A16 strategies results in negligible loss compared to the original model, while using the W4A16 strategy leads to significant loss in out-of-domain scenarios. Regarding pruning, all pruned models exhibit substantial losses in out-of-domain scenarios, yet own relatively minor loss in in-domain scenarios.